For the next fiscal year, you forecast net income of $50,000 and ending assests of $500,000. Your firm payout ratio is 10%. Your beginning stockholders equity is $300,000 and your beginning total liabilities are $120,000. Your non-debt liabilites such as accounts payable are forecasted to increase by $10,000. Assume your beginning debt is $ 100000. What amount of equity and what amount of debt would you need to issue to cover the net new financing in order to keep your​ debt-equity ratio​ constant?
1
Expert's answer
2018-02-13T09:17:07-0500
1) Beginning debt-equity ratio constant =beginning stockholders equity / beginning debt = $300000/($100000+$120000) = 1.36
2) Final debt-equity ratio constant = final stockholders equity / final debt = 1.36
Retained earnings = $50000 – ($50000*0.1) = $45000
final stockholders equity = $269067.8+$45000= $314067.8
Final debt-equity ratio constant = $314067.8/($500000 - $269067.8) = 1.36